# Auto-Generating Earnings Report Analysis via an Augmented LLM

### v1

@@ Abstract
Financial analysis heavily relies on the evaluation of earnings reports to gain insights into company performance. Traditional generation of these reports requires extensive financial expertise and is time-consuming. With the impressive progress in Large Language Models (LLMs), a wide variety of financially focused LLMs has emerged, addressing tasks like sentiment analysis and entity recognition in the financial domain. This paper presents a novel challenge: developing an LLM specifically for automating the generation of earnings reports analysis. Our methodology involves an in-depth analysis of existing earnings reports followed by a unique approach to fine-tune an LLM for this purpose. This approach combines retrieval augmentation and the generation of instruction-based data, specifically tailored for the financial sector, to enhance the LLM’s performance. With extensive financial documents, we construct financial instruction data, enabling the refined adaptation of our LLM to financial contexts. Preliminary results indicate that our augmented LLM outperforms general open-source models and rivals commercial counterparts like GPT-3.5 in financial applications. Our research paves the way for streamlined and insightful automation in financial report generation, marking a significant stride in the field of financial analysis.
Keywords:
LLMs financial analysis instruction tuning.
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable proficiency in diverse applications, including financial analysis. BloombergGPT [7], a specialized LLM for finance, is built from scratch and excels in various financial tasks, though it is costly due to its use of proprietary data and training expenses of over \$2 million. Conversely, FinGPT [8] is a low-cost, open-source LLM fine-tuned on instructional data, focusing on basic financial tasks like sentiment classification and named entity recognition. This paper explores more complex financial challenges, particularly earnings report analysis.
@@
In addressing domain-specific challenges, an LLM frequently leverages the retrieval augmented generation (RAG) approach [3], a methodology that furnishes the model with domain-specific context. In our work, aimed at grounding generation instructions within the financial domain, we introduce a novel method termed retrieval-augmented instruction data generation. This methodology integrates both the financial domain context and the instructional prompt during data collection, guiding the teacher LLM in generating questions and answers aligned with the provided context. The approach yields a rich set of financial instruction-following data, important for the fine-tuning of a financial-augmented LLM.
@@
In summary, this research makes several significant contributions. Firstly, we address the earnings report analysis problem, representing a pivotal extension in the realm of LLMs focused on finance. Our novel approach involves a retrieval-augmented instruction data generation method tailored for domain-specific tasks. This method generates instruction data contextually grounded in the financial domain. By fine-tuning an LLM on the curated instruction data, we demonstrate superior performance within the financial domain, much better than the well-known open-source LLM, Llama-2-7b [6], and comparable with the commercial GPT-3.5 model.
2 Earnings Report Analysis
Properly analyzing an earnings report of a company is a crucial task for any analyst because it involves a comprehensive understanding of the company’s financial health and performance. A few guidelines are provided on the web as to how to perform the analysis, but there is no definite standard on what should be included in the earnings report. Moreover, the analysis involves not just an objective evaluation of the numbers but also a subjective choice of which financial numbers to look for from a company’s earnings reports. Those components are what we call the Key Performance Indicators (KPIs). There are many components that make up the earnings analysis, but KPIs are one of the most basic and important numbers that the analyst chooses.
The KPIs are what we will use as the main content for auto-generating the analysis report. It is important to note that no KPIs are the same for each company and for each analyst. During the earnings call, the CEO usually chooses some set of numbers to argue that the company is doing great. But, the problem is that those numbers are chosen differently over time. Additionally, from the analyst’s point of view, he may not look for those numbers but rather needs the numbers from his own set of KPIs. These together raise discrepancies in the company’s analysis.
To address that discrepancy, we introduce a number of baseline KPIs that the analyst can use to initially gauge the company’s financial performance. Then, the analyst can manually update them to customize the analysis. The report will then be generated based on the results of those KPIs. In other words, it involves human-in-the-loop feedback from the analyst, which then prompts the model to auto-generate the analysis report. Adopting KPIs into our report generation process preserves the uniqueness of the analyst’s point of view while maintaining the consistency of the process. In the experiment part, we evaluate our model with basic KPI-based questions like the quarterly dividend.
3 Retrieval Augmented Instruction Tuning
This section describes our retrieval-augmented instruction tuning method and how to build a financially augmented LLM by fine-tuning from general financial instruction data.
@@ 3.1 General financial instruction data
Table 1: Instruction prompt to generate general financial instruction data and a data sample of generated instruction-following data. $\{num\_questions\_per\_chunk\}$ is the number of data samples we want to generate for a chunk of text.
Instruction: You are an earnings report analyst. Your task is to ask $\{num\_questions\_per\_chunk\}$ questions to understand a company, its financial report, and its key financial performance. The questions should be diverse in nature across the document. Restrict the questions to the context of the information provided.
@@
To fine-tune a financially augmented LLM, we need general financial instruction data that can be generated from the context of financial documents, such as quarterly reports. This process entails initially segmenting the financial document into discrete text chunks. Subsequently, each text chunk serves as the context for the generation of instruction-following data pertaining to financial nuances. Employing a teacher LLM, such as GPT-3.5-turbo, we extract insights and generate relevant instructions within the financial context. The specifics of this process, along with an illustrative example, are delineated in table 1. We opt for a configuration with ${num\_questions\_per\_chunk}$ set to 10, amassing a dataset of 400 financially augmented instruction instances derived from a single financial report.
Figure 1: The retrieval augmented instruction data generation workflow.
In Fig. 1, we present the workflow of our innovative retrieval-augmented instruction data generation methodology. Commencing with a domain-specific document indexing step, we populate a vector database (such as ChromaDB) to establish a contextual foundation within the domain. Subsequently, the instruction prompt undergoes processing by the teacher LLM (such as GPT-3.5), which generates pertinent question-answer pairs based on the extracted contextual information. This comprehensive approach ensures that the instruction generation is intricately tied to the relevant domain context, enhancing the quality and specificity of the resultant data.
@@ 2 Key financial indicators What is the revenue? Press Release, Earnings Report
Precise guidance for the teacher LLM in the form of seed instructions is essential to ensure the generation of accurate and varied instruction-following data. Given the intricacies involved in earnings analysis, we employ conversation-based instruction prompts tailored to facilitate the analysis of a company over a specific period. This contextual awareness enables the model to comprehend the entire analysis methodology, contributing to its ability to generate coherent and insightful earnings analysis. A comprehensive breakdown of various instruction types, accompanied by illustrative prompt examples and their corresponding financial documents, is presented in table 2.
3.3 Augmented LLM
Our primary goal is to harness the capabilities of a pre-trained LLM in conjunction with financially augmented instruction data. In this work, we adopt Llama-2-7b [6] as the foundational pre-trained LLM and employ LoRA [2] method for efficient parameter-specific fine-tuning. To enhance memory efficiency and expedite the training process, we incorporate a 4-bit compression technique following the QLoRA method [1]. The training data has the following format, in which $sample$ is a training sample of the financial instruction dataset.
We have provided context information below.
{sample[‘context’]}
Given this information, please answer the question:
{sample[‘query’]}
Answer: {sample[‘answer’]}
@@ 4 Experimental results
4.1 Experiment set up
We employed the Google Cloud virtual machine for training on four Tesla T4 GPUs, each equipped with 16GB of RAM. The training process spanned three epochs, maintaining a consistent learning rate of 2e-4 while adhering to a maximum sequence length of 2048 for optimal model performance. The total training time was 40 hours.

### v3

# RAG-IT: Retrieval-Augmented Instruction Tuning for Automated Financial Analysis - A Case Study for the Semiconductor Sector

@@ Abstract
Financial analysis relies heavily on the interpretation of earnings reports to assess company performance and guide decision-making. Traditional methods for generating such analyzes require significant financial expertise and are often time-consuming. With the rapid advancement of Large Language Models (LLMs), domain-specific adaptations have emerged for financial tasks such as sentiment analysis and entity recognition. This paper introduces RAG-IT (Retrieval-Augmented Instruction Tuning), a novel framework designed to automate the generation of earnings report analysis through an LLM fine-tuned specifically for the financial domain. Our approach integrates retrieval augmentation with instruction-based fine-tuning to enhance factual accuracy, contextual relevance, and domain adaptability. We construct a sector-specific financial instruction dataset derived from semiconductor industry documents to guide the LLM adaptation to specialized financial reasoning. Using NVIDIA, AMD, and Broadcom as representative companies, our case study demonstrates that RAG-IT substantially improves a general-purpose open-source LLM and achieves performance comparable to commercial systems like GPT-3.5 on financial report generation tasks. This research highlights the potential of retrieval-augmented instruction tuning to streamline and elevate financial analysis automation, advancing the broader field of intelligent financial reporting.
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable proficiency in diverse applications, including financial analysis. BloombergGPT [15], a specialized LLM for finance, is built from scratch and excels in various financial tasks, though it is costly due to its use of proprietary data and training expenses of over \$2 million. Conversely, FinGPT [16] is an open-source low-cost LLM fine-tuned on instructional data, focusing on basic financial tasks like sentiment classification and named entity recognition. This paper explores more complex financial challenges, particularly earnings report analysis.
@@
In addressing domain-specific challenges, an LLM frequently leverages the retrieval augmented generation (RAG) approach [8], a methodology that provides the model with domain-specific context. In our work, aimed at grounding generation instructions within the financial domain, we introduce a novel method termed retrieval-augmented instruction data generation. This methodology integrates both the financial domain context and the instructional prompt during data collection, guiding the teacher LLM in generating questions and answers aligned with the provided context. The approach yields a rich set of financial instruction-following data, important for the fine-tuning of a financial-augmented LLM.
@@
In summary, this research makes several significant contributions. Firstly, we address the earnings report analysis problem, representing a pivotal extension in the realm of LLMs focused on finance. Our novel approach involves a retrieval-augmented instruction data generation method designed for domain-specific tasks. This method generates instruction data contextually grounded in the financial domain. By fine-tuning an LLM with the curated instruction data, we demonstrate superior performance within the financial domain, much better than the well-known open-source LLM, Llama-2-7b [12], and comparable with the commercial GPT-3.5 model.
2 Structured Components of Earnings Report Analysis
A comprehensive earnings report analysis goes beyond the computation of financial metrics; it involves a structured reasoning process across several interconnected dimensions of company performance. To ensure both interpretability and reproducibility of the analysis, we decompose the process into a series of analysis modules, each corresponding to a specific section of an earning report. These modules form the basis of our data-generation and report-synthesis framework.
Each instruction type encapsulates a distinct analytical objective, as detailed below. To ensure that our six-component structure is not arbitrary, we consulted two financial experts with MBA degrees and many years of experience in large corporations. Their feedback confirmed that this decomposition aligns well with the workflow used in professional earnings reports, providing practical grounding for our design choices.
(1) Company Core Information
This section introduces the company’s background, business model, and sectoral context. It provides foundational context for the subsequent financial and strategic assessments. Typical questions at this level include identifying the company’s primary business domains, product lines, and target markets. The main information sources are the company’s press releases and official websites, which ensure factual accuracy. This step grounds the report-generation process, enabling the model to contextualize later KPI-based interpretations within the company’s operational scope.
(2) Key Financial Indicators (KPIs)
The next stage involves extracting and interpreting core quantitative metrics that define the company’s financial performance. Examples include revenue, net income, operating margin, earnings per share (EPS), and cash flow. These KPIs are retrieved from earnings reports and press releases, then aligned with the analyst’s custom-defined KPI set. This allows the analysis to be both standardized and flexible-standardized in structure, yet personalized to reflect the analyst’s expertise. In our framework, these KPIs serve as anchor points for the model’s narrative generation, allowing the auto-generated analysis to maintain factual grounding while adapting to subjective analytical focus. Indeed, financial KPIs have been characterized as metrics that provide insights into underlying business performance across categories such as profitability, liquidity, efficiency, valuation and leverage. [13, 1]
(3) Comparative Analysis
To provide market context, comparative analysis benchmarks the company’s performance against peers in the same industry. This may involve ratio-based comparisons (e.g., revenue growth rate, profit margin) or relative ranking within the sector. The comparison step enhances interpretability by situating a company’s performance within broader market trends. The relevant documents include both earnings reports of competitors and industry summaries, allowing the model to contextualize differences in business cycles or strategic focus.
(4) Outlook Analysis
Forward-looking analysis assesses management’s guidance, analyst consensus, and projected financial trends. By drawing from press releases, earnings reports, and occasionally analyst forecasts, the model can synthesize expectations for future performance, such as projected revenue growth or cost-optimization strategies. This component also reflects qualitative signals, such as management tone or sentiment extracted from the earnings call transcript, which complement the quantitative KPIs. Integrating these elements allows the model to generate a balanced view that combines data-driven projections with managerial perspectives.
(5) Summary and Executive Sentiment
The summary component synthesizes key insights from both financial and textual sources, such as executives’ remarks and press announcements. This step aims to distill the overall narrative conveyed during the earnings call, highlighting management’s confidence levels, strategic priorities, or risk acknowledgments. We use this component to train the model to capture soft indicators that are not directly observable in numerical data but are critical for evaluating corporate outlook and investor sentiment.
(6) Integrated Analysis Report
Finally, all the preceding elements are integrated into a cohesive earnings report analysis. This module combines objective KPI evaluation, comparative benchmarks, and qualitative outlook synthesis to generate a holistic view of the company’s performance in the given quarter. The structure and tone are inspired by professional equity research reports, ensuring the generated output is aligned with financial reporting standards. The analyst remains in the loop by reviewing or updating the KPI set, thereby guiding the model’s focus in subsequent iterations.
3 Financial Instruction Data
Developing a financially augmented large language model (LLM) requires a well-structured and context-rich dataset that reflects the language, reasoning patterns, and analytical conventions of financial reporting. To achieve this, we construct a two-tier dataset which serves as the foundation for instruction tuning and retrieval-augmented fine-tuning.
@@ 3.1 General Financial Instruction Data
Table 1: Instruction prompt to generate general financial instruction data and a data sample of generated instruction-following data. num_questions_per_chunk{num\_questions\_per\_chunk} is the number of data samples we want to generate for a chunk of text.
Instruction: You are an earnings report analyst. Your task is to ask num_questions_per_chunk{num\_questions\_per\_chunk} questions to understand a company, its financial report, and its key financial performance. The questions should be diverse in nature across the document. Restrict the questions to the context of the information provided.
@@
To fine-tune a financially augmented LLM, we first require a diverse set of general financial instruction data that reflect real-world financial document structures and semantics. Such data are generated from the contextual content of company financial disclosures, including quarterly and annual reports, management commentary, and earnings call summaries.
The data generation process begins by segmenting each financial document into discrete, semantically coherent text chunks. Each chunk represents a self-contained financial context that can independently support the generation of multiple instruction-following pairs. Examples of such contexts include a company’s identity, its industry classification, and reported key financial metrics. For instance, our segmentation of NVIDIA’s quarterly earnings report yielded chunks addressing queries such as:
  • What is NVIDIA?
  • What are the main industry sectors of NVIDIA?
  • What is the revenue of NVIDIA for the second quarter of fiscal year 2024?
  • What is the GAAP and Non-GAAP earnings per diluted share for the quarter?
  • How many shares did NVIDIA repurchase during the second quarter of fiscal 2024, and for what amount?
Each chunk thus acts as a contextual unit from which the model can learn to map document-grounded financial text to relevant natural language queries. Using a teacher LLM such as GPT-3.5-turbo, we automatically generate instruction–response pairs based on the textual content of each chunk. The teacher model is prompted to extract domain-specific insights, numerical facts, and relational patterns (e.g., quarter-over-quarter growth, year-over-year comparison, dividend schedules, or share repurchases).
To ensure balanced coverage, we set the configuration parameter num_questions_per_chunk{num\_questions\_per\_chunk} to 10. For a single quarterly report containing approximately 40 well-defined chunks, this setup yields around 400 financial instruction–response pairs. These pairs collectively form the general financial instruction dataset, which serves as the base layer for fine-tuning the financially augmented LLM.
The resulting dataset encompasses multiple dimensions of financial understanding, including:
  1. 1.
    Entity-level understanding: Identifying the company, fiscal period, and report structure.
  2. 2.
    Numerical reasoning: Computing and interpreting KPIs such as revenue, operating income, and EPS.
  3. 3.
    Temporal reasoning: Comparing performance across quarters and fiscal years.
  4. 4.
    Policy and governance insight: Understanding board authorizations, dividend policies, and share repurchase programs.
This structured data generation process enables the LLM to capture the linguistic and analytical characteristics of financial narratives, providing a strong foundation for subsequent retrieval-augmented fine-tuning. An illustrative sample of the generated instruction-following data is shown in Table 1.
In summary, the general financial instruction dataset serves two purposes: (1) it provides a systematic foundation of financial question–answer pairs derived directly from authentic corporate disclosures, and (2) it establishes the contextual grounding necessary for extending the LLM’s capacity to handle retrieval-augmented financial reasoning tasks discussed in the next section.
@@ 2 Key financial indicators What is the revenue? Press Release, Earnings Report
Precise guidance for the teacher LLM in the form of seed instructions is essential to ensure the generation of accurate, diverse, and contextually grounded instruction-following data. Given the complexities inherent in financial text, particularly in earnings reports, the model must understand not only numerical facts but also qualitative relationships, comparative metrics, and managerial sentiment. To this end, we design a structured set of seed instructions that guide the teacher model to produce high-quality financial reasoning data.
The seed instructions are crafted to emulate the reasoning process of human financial analysts. They cover multiple analytical dimensions, including factual extraction, comparison, temporal reasoning, and interpretive commentary. We employ conversation-based instruction prompts to create a natural flow of inquiry-mimicking how an analyst might sequentially explore a company’s quarterly performance. This contextual chaining enables the teacher model to capture dependencies between questions and answers, resulting in coherent and analytically rich datasets.
Table 2 presents a taxonomy of six instruction types with representative examples and their associated document sources. Each instruction type plays a distinct role in the generation of financial instruction-following data:
  • Company Core Information: Focuses on the company’s identity, business model, and operational domain. These questions establish the foundational context of the analysis.
  • Key Financial Indicators: Targets quantitative performance metrics such as revenue, earnings per share (EPS), or operating margin. The responses typically include both the reported values and brief interpretations.
  • Comparison: Involves benchmarking company metrics against peer firms or historical performance to capture relative strengths or weaknesses.
  • Outlook: Encapsulates management guidance, analyst consensus, and forward-looking statements that indicate projected financial performance. These prompts encourage narrative synthesis beyond raw numbers.
  • Summary: Integrates insights from multiple sections of the earnings release and call transcripts, summarizing the quarter’s highlights and strategic focus.
  • Analysis: Represents the highest level of reasoning, combining numerical interpretation, strategic evaluation, and investor sentiment to form a holistic analysis. This aligns most closely with the structure of an equity research report.
Each instruction type is linked to specific source documents—such as press releases, earnings reports, call transcripts, or equity research summaries—to ensure grounding in verified financial text. During instruction data generation, the teacher LLM is conditioned on both the prompt and its associated document snippet, allowing the resulting responses to be evidence-based and contextually aligned.
Furthermore, we design the prompting sequence to include interdependent instructions, where later prompts (e.g., Outlook or Analysis) depend on information extracted in earlier steps (e.g., Key Financial Indicators). This multi-turn structure improves the model’s capacity to integrate cross-referenced information, mirroring the analytical reasoning process of human experts.
The seed instructions thus serve as a blueprint for generating a comprehensive and multi-dimensional financial dataset. By training the student model on these structured instruction–response pairs, the LLM learns to:
  1. 1.
    Accurately extract and interpret quantitative metrics from textual disclosures.
  2. 2.
    Contextualize these metrics within industry and temporal frames.
  3. 3.
    Generate coherent narratives that explain the causes and implications of financial outcomes.
In summary, earnings analysis seed instructions establish a consistent and interpretable framework for generating high-quality financial reasoning data. They ensure that the model learns not only to retrieve factual content but also to perform structured analysis—an essential step toward developing a retrieval-augmented financial LLM capable of producing explainable and analyst-like insights.
4 Retrieval-Augmented Instruction Tuning Method
This section describes our retrieval-augmented instruction tuning method and outlines how we construct a financially augmented large language model (LLM) through fine-tuning using general financial instruction data and retrieval-augmented financial reasoning data. The overall framework aims to enhance an LLM’s ability to reason about financial information in context, improving its performance on document-grounded financial tasks.
Figure 1: The retrieval-augmented instruction data generation workflow.
In Fig. 1, we present the workflow of our retrieval-augmented instruction data generation and model fine-tuning methodology. The approach is designed to tightly couple context retrieval, instruction generation, and parameter-efficient fine-tuning. The method consists of two main components: (1) the retrieval-augmented instruction data generation process, and (2) the financially augmented LLM fine-tuning process.
4.1 Retrieval-Augmented Instruction Data Generation
This component focuses on creating context-grounded, instruction-following financial data derived from authentic financial documents. The goal is to generate diverse and accurate question–answer pairs that reflect the reasoning tasks encountered in financial analysis.
The pipeline begins with a document indexing and retrieval stage. Financial documents such as quarterly earnings reports, press releases, and SEC filings are segmented into coherent text chunks. These chunks are then embedded using a dense embedding model and stored in a vector database (e.g., ChromaDB), which supports semantic retrieval.
Given a financial topic or company name, the system retrieves relevant chunks based on similarity search. These chunks are passed as contextual input to a teacher model (e.g., GPT-3.5-turbo), which generates instruction–response pairs grounded in the retrieved financial context. Each generated sample follows the structure:
  "instruction": <financial query>,
  "input": <retrieved document context>,
  "output": <teacher-generated answer>
This process ensures that generated data remain faithful to real financial statements and reduce hallucinations by grounding responses in retrieved evidence. In addition, conversational-style prompting is employed to elicit richer, more analytical responses that capture patterns such as revenue growth, expense variation, or year-over-year comparisons.
After generation, the data undergo normalization (e.g., numeric unit alignment, temporal standardization, duplicate removal). The final dataset integrates both general and analytic instructions, supporting a wide range of reasoning patterns that align with financial interpretation and decision-making.
4.2 Financially Augmented LLM Fine-Tuning
Once the retrieval-augmented financial instruction dataset is prepared, we fine-tune a pre-trained LLM to develop a financially specialized reasoning capability. The fine-tuning procedure focuses on efficiency and stability while preserving general language understanding.
During training, each instruction sample (samplesample) follows the previously described schema, where the instruction represents a financial query, the input provides relevant context, and the output serves as the correct answer. The model learns to interpret financial context, retrieve relevant details, and synthesize grounded analytical conclusions.
This approach ensures that the fine-tuned LLM can:
  • Perform context-grounded financial question answering with factual precision.
  • Conduct multi-hop reasoning, such as temporal comparisons or cross-metric correlations.
  • Generate human-like financial commentary aligned with analyst-style interpretations.
In summary, the retrieval-augmented instruction tuning framework integrates contextual data generation and parameter-efficient fine-tuning to build a domain-adapted financial LLM. The retrieval process ensures factual grounding, while instruction tuning refines the model’s analytical expressiveness. Together, these two components yield a system capable of delivering accurate, interpretable, and contextually grounded financial reasoning.
@@ 5 Experimental results
5.1 Experiment configurations and set up
We employ Llama-2-7B [12] as the base model and fine-tune it using the Low-Rank Adaptation (LoRA) method [6]. Following the QLoRA setup [2], we enable 4-bit (NF4) quantization using prepare_model_for_kbit_training. Our LoRA configuration uses a rank of 64, lora_alpha = 16, and a dropout of 0.1, applied to all attention and feed-forward projection layers with no bias adaptation. Fine-tuning is conducted using SFTTrainer with a maximum sequence length of 1024 tokens and packing enabled to improve training throughput.
We train for 3 epochs with a global batch size of 8 (per-device batch size 2, gradient accumulation 2) using the paged_adamw_32bit optimizer. The learning rate is set to 2e-5 with a constant schedule and a warmup ratio of 0.03. We disable gradient checkpointing and use ddp_find_unused_parameters=False for efficient multi-GPU execution. The maximum gradient norm is clipped to 0.3.
We employed the Google Cloud virtual machine for training on four Tesla T4 GPUs, each equipped with 16GB of RAM. Due to the limited compute capability and memory bandwidth of T4 GPUs combined with 4-bit QLoRA dequantization overhead and activation packing, the end-to-end fine-tuning of our 800-sample dataset requires approximately 40 hours.
For the retrieval component, we use the text-embedding-3-large encoder to embed documents. The corpus is chunked into 512-token segments with a 128-token overlap. At inference time, the retriever selects the top k=3 most relevant chunks using cosine similarity.
To ensure the correctness of GPT-3.5–generated training data, we apply a two-stage quality control process: (1) rule-based filtering (format validation, length constraints, and inconsistency checks), and (2) manual inspection of a randomly sampled 15% subset to identify hallucinations, numerical inconsistencies, and unclear instructions. Problematic items are revised or removed before final training.
@@ 5.2 Evaluation metrics
The prompting for GPT-4 judgment is show in the listing 1.
Listing 1: GPT-4 Judge Prompt
You are an expert evaluator. Your task is to compare a candidate answer
against a reference answer and assign a numeric score based on the
scoring rubric provided. You must follow the rubric strictly and output
only the final score as an integer from 1 to 10.
Score 1: The answer is completely unrelated to the reference.
Score 3: The answer has minor relevance but does not align with the
reference.
Score 5: The answer has moderate relevance but contains inaccuracies.
Score 7: The answer aligns with the reference but has minor errors.
Score 9: The numbers match perfectly, structure is similar.
Score 10: Completely accurate and perfectly aligned.